OpenStudy (anonymous):

Hi does anybody mind helping w. 4 questions? Willing to give fan and medal! I guess? :)

OpenStudy (anonymous):

Jake has proved that a function, f(x), is a geometric sequence. How did he prove that? He showed that an explicit formula could be created. He showed that a recursive formula could be created. He showed that f(n) ÷ f(n - 1) was a constant ratio. He showed that f(n) - f(n - 1) was a constant difference.

OpenStudy (anonymous):

A store had 250 bottles of water. Each week, 40% of the bottles were sold and 48 new bottles arrived in shipments. Which recursive function best represents the number of bottles of water in the store, given that f(0) = 250? f(n) = f(n - 1) • 0.6 + 48, n > 0 f(n) = 250 - f(n - 1) • 0.4 + 48, n > 0 f(n) = f(n - 1) • 0.4 + 48, n > 0 f(n) = 250 - f(n - 1) • 0.6 + 48, n > 0

OpenStudy (anonymous):

Barry has been watching the geese that live in his neighborhood. The number of geese changes each week. n f(n) 1 56 2 28 3 14 4 7 Which function best shows the relationship between n and f(n)? f(n) = 28(0.5)n f(n) = 56(0.5)n-1 f(n) = 56(0.5)n f(n) = 112(0.5)n-1